Let v and w be two vectors whose modules are equal to 3 and 1, respectively the angle formed between them is equal to π / 6. If θ denotes the measure of the angle between v + w and v-w, how much is cos θ worth?
Answer #1

Given v=3 IW =1 let angle between Vectors vzw is a idatura Tofind [v+) FJV-WI Gangle between vetos is o [v+w= mit rot + Rivol|(146) +(-6)*= twitw-witowv+w/-w9) Coso 12V1? = W 10t313 )4J10-20 )4 210t. 35/J10-36)COLA 4MP = 1043/3 410-83+ 2 J109-LB3) c